OpenStudy (anonymous):

solve for x from [0,2pi): sqrt3 tanx=1

OpenStudy (anonymous):

\[\sqrt{3}\tan x=1\]

OpenStudy (anonymous):

divide by sqrt3 both sides: tanx = 1/sqrt3 = (sqrt3)/3 x is a common angle... [HINT]

OpenStudy (anonymous):

its 30 degrees right?

OpenStudy (anonymous):

yes.. but there's another one... btw... they asked for radians so 30 degrees = pi/6

OpenStudy (anonymous):

it's in quadrant 3 that's where tan is positive.

OpenStudy (anonymous):

7pi/6?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

good job kaisan... :) btw... maybe you should write out the tangents of each of those common angles on your unit circle... ... just a suggestion.

OpenStudy (anonymous):

i'm not really good with tangent, i always get things mixed up

OpenStudy (anonymous):

and that's why u should write it on your unit circle...:)